OpenStudy (anonymous):

Can you guys help me with something real quick? I'm a bit confused with this math homework. I'll give medals.

OpenStudy (anonymous):

you can only give one meda. but shoot.

OpenStudy (anonymous):

Imagine that a generous relative has agreed to give you $3 for every $2 you save. Use a part whole model to calculate how much of the savings goal will be yours to raise and how much you can expect the relative to contribute. The numbers I have for the trip are: Actual cost: 8,100 Rounded cost: 8,800 I'm really confused on this.

OpenStudy (anonymous):

spend $2 ------> make $3 now, what is this about the actual and total cost?

OpenStudy (anonymous):

How much money would I need raise and how much will the relative provide for the trip?

OpenStudy (anonymous):

i see ok lets say you raise "x" then your relative should pay -> \(\Large {x\times 3\over2}\)

OpenStudy (anonymous):

so, total expenditure = money you put in + provided by relative total = x + \(\Large{x\times3\over2}\)

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

OK.. I'm still confused.

OpenStudy (anonymous):

let us start with the assumption that we put in "x" amount

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

then the relative puts in a proportional amount: for every "$2" that you put in, the relative contributes "$3" so, for "$x" that you put in, the relative contributes ______

OpenStudy (anonymous):

you see it? the relative contributes \[x\times 3\over2\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

gud. then the total amount you've got: = the amount you put in + the amount from the relative

OpenStudy (anonymous):

how much amount do you need in total?

OpenStudy (anonymous):

So it's 9,000 x 3?

OpenStudy (anonymous):

no wait.. we need to know how much we need in total

OpenStudy (anonymous):

how much does the trip cost?

OpenStudy (anonymous):

Well for the whole 3 trips I've 'planned' cost 5,500, 1,500, 1,100

OpenStudy (anonymous):

i see so you will need to raise a total of = 5500 + 1500 + 1100 = 8100 correct?

OpenStudy (anonymous):

Wait the first one is 4,000 sorry. :\

OpenStudy (anonymous):

ok, so how much do you need to raise in total then?

OpenStudy (anonymous):

so it would be 4,000 + 1,500 +1,100 = 6,600 \

OpenStudy (anonymous):

good. this amount = amount you put in + amount relative puts in \[6600=x+{x\times3\over2}\] solve for "x"

OpenStudy (anonymous):

x is 3?

OpenStudy (anonymous):

what? no.. you have to solve for it.

OpenStudy (anonymous):

So to solve it I have to divide right?

OpenStudy (anonymous):